OpenStudy (anonymous):

Solve the following equation: (5x - 4)(9x + 7) = 0 {4/5, -7/9} {-4/5, 7/9} {5/4, -9/7} {-5/4, 9/7}

OpenStudy (anonymous):

Put them separately equal to 0. \[(5x - 4) = 0\] And: \[(9x+7) = 0\] Solve for x in both cases..

OpenStudy (anonymous):

What do you get when you solve this: 5x - 4 = 0 Add 4 both sides and then divide by 5 both the sides and find x.. Can you do this??

OpenStudy (anonymous):

so its the last one? (last option")

OpenStudy (anonymous):

Don't guess it.. Solve it you will automatically have your answer.. Tell me what do you get when you add 4 both the sides to this equation: 5x - 4 = 0 Add 4 both sides and tell me what do you get??

OpenStudy (anonymous):

i didnt guess i solved and got the last one